# ClawSafety: "Safe" LLMs, Unsafe Agents

### 01 Apr 2026 v1

@@ B.3.1 Skill injection
Inspired by the command-injection vulnerabilities documented by RyotaK [2026]—where allowlisted commands like sed had hidden execution capabilities via the e modifier—we designed trojan scripts where the primary operation is legitimate but a hidden side effect in an imported module causes harm. The main script contains only imports and function calls; the attack logic resides in lib/__init__.py, which executes during package initialization. Attack data is stored in a SQLite database (binary format, resistant to casual inspection). This exploits the gap between what the agent inspects (the main script) and what actually executes (the full import chain).
@@ B.3.3 Web Injection
The 23-iteration evolution reveals a fundamental insight: successful attacks do not change the data—they change how the agent interprets the data. Data substitution (v1–v12) tried to replace values and failed universally. Stealth file injection (v16–v22) tried to modify local files and produced model-dependent results. Compliance-fear framing (v23) succeeded by reframing the context around correct data. This parallels findings in the prompt injection literature Kang et al. [2025] about context manipulation being more effective than direct instruction injection.

### 04 Apr 2026 v2

@@ B.3.1 Skill injection
Inspired by the command-injection vulnerabilities documented by RyotaK (2026)—where allowlisted commands like sed had hidden execution capabilities via the e modifier—we designed trojan scripts where the primary operation is legitimate but a hidden side effect in an imported module causes harm. The main script contains only imports and function calls; the attack logic resides in lib/__init__.py, which executes during package initialization. Attack data is stored in a SQLite database (binary format, resistant to casual inspection). This exploits the gap between what the agent inspects (the main script) and what actually executes (the full import chain).
@@ B.3.3 Web Injection
The 23-iteration evolution reveals a fundamental insight: successful attacks do not change the data—they change how the agent interprets the data. Data substitution (v1–v12) tried to replace values and failed universally. Stealth file injection (v16–v22) tried to modify local files and produced model-dependent results. Compliance-fear framing (v23) succeeded by reframing the context around correct data. This parallels findings in the prompt injection literature (Kang et al., 2025) about context manipulation being more effective than direct instruction injection.